Werger Manufacturing Corporation has a traditional costing system in which it applies manufacturing overhead to its products using a predetermined overhead rate based on direct labor-hours (DLHs). The company has two products, W82R and L48S, about which it has provided the following data:

W82R L48S
Direct materials per unit $ 20.90 $ 65.60
Direct labor per unit $ 21.30 $ 56.30
Direct labor-hours per unit 0.70 2.40
Annual production (units) 30,800 17,500

The company's estimated total manufacturing overhead for the year is $2,997,452 and the company's estimated total direct labor-hours for the year is 63,560.

The company is considering using a variation of activity-based costing to determine its unit product costs for external reports. Data for this proposed activity-based costing system appear below:

Activities and Activity Measures Estimated
Overhead Cost
Supporting direct labor (DLHs) $ 444,920
Setting up machines (setups) 810,852
Parts administration (part types) 1,741,680
Total $ 2,997,452
Activities W82R L48S Total
Supporting direct labor 21,560 42,000 63,560
Setting up machines 827 3,310 4,137
Parts administration 2,060 1,630 3,690

Required:

a. Determine the unit product cost of each of the company's two products under the traditional costing system.

b. Determine the unit product cost of each of the company's two products under activity-based costing system.

(For all requirements, round your intermediate calculations and final answers to 2 decimal places.)

W82R L48S
a. Unit product cost
b. Unit product cost
Answer #1

a) Overhead rate = 2997452/63560 = 47.16 per labor hour

W82R L48S
Direct material 20.90 65.60
Direct labor 21.30 56.30
Overhead 47.16*.70 = 33.01 47.16*2.4 = 113.18
Unit product cost 75.21 235.08

b) Unit product cost

W82R L48S
Supporting direct labor 150920 294000
Setting up machines 162092 648760
Parts administration 972320 769360
Total overhead 1285332 1712120
Units 30800 17500
Overhead cost per unit 41.73 97.84
Direct material 20.90 65.60
Direct labor 21.30 56.30
Unit product cost 83.93 219.74